OpenStudy (anonymous):

I cant get the recurrence relation to work at all, could someone give a brief summary on how to do 2y''+4xy=0 expansion is sumation A_n[x-x_0]^n centered at x_0= 0 i have absolutely no idea how to do this!

OpenStudy (amistre64):

let y = a0 + a1x + a2 x^2 + a3 x^3 + .... y'' = 2 a2 + 6 a3 x + 12 x^2 + ... or written in summations:\[y=\sum_0a_n~x^n\] \[y''=\sum_2a_n~n(n-1)x^{n-2}\] plugging this in we get: \[2\sum_2a_n~n(n-1)x^{n-2}+4x\sum_0a_n~x^n=0\] and smoothing things thru \[\sum_2~2a_n~n(n-1)x^{n-2}+\sum_04a_n~x^{n+1}=0\] adjust the indexes/exponent to the lowest form \[\sum_2~2a_n~n(n-1)x^{n-2}+\sum_{0+3}4a_{n-3}~x^{n-3+1}=0\] \[\sum_2~2a_n~n(n-1)x^{n-2}+\sum_{3}4a_{n-3}~x^{n-2}=0\] and pop out a term to line up the summation indexes \[4a_2+\sum_3~2a_n~n(n-1)x^{n-2}+\sum_{3}4a_{n-3}~x^{n-2}=0\] and combine the summations \[4a_2+\sum_{3}2a_n~n(n-1)+4a_{n-3}~x^{n-2}=0\] from this, we can determine that a2=0, and the coefficients need to go to zero as well \[2a_n~n(n-1)+4a_{n-3}=0\] solve for an \[a_n=-\frac{2a_{n-3}}{n(n-1)}~:~n>2\] start defining the terms \[a_0=a_0, a_1=a_1, a_2=0\] \[a_3=-\frac{2a_{0}}{3(2)}\] \[a_4=-\frac{2a_{1}}{4(3)}\] \[a_5=-\frac{2a_{2}}{5(4)}\] \[a_6=-\frac{2a_{3}}{6(5)}\] since all the a2 parts, and their related coeffs are zero, we can ignore those and define the rest in terms of a0 and a1